OpenStudy (anonymous):

Class has 50 question test, each correct answer is 2 pt, each wrong is -.5, if no answer then no points. If student needs to earn 70 pt on the test then what is the equation in terms of incorrect answers

OpenStudy (triciaal):

number correct = x number incorrect = 50-x when all correct x = 50 ; points = 100 when all incorrect 50*(-0.5) = -25 points to get 70 points 2x -0.5(50-x)= 70

OpenStudy (anonymous):

so you dont solve for x but state its 50?

OpenStudy (triciaal):